# Reconstruction of Differentially Private Text Sanitization via Large Language Models

### v2

@@ Appendix E Large Language Models Used in the Experiments
Appendix F Ablation Studies
The original text of the DP-santized one was not seen by LLMs. Table 16 shows the results of prompting fine-tuned white-box LLMs with DP-sanitized text, where the original text was not seen by the LLMs.
Privacy budgets. In this experiment, we aim to investigate how privacy budget affects the quality of reconstruction attacks. Intuitively, privacy budgets control the level of modification for the DP text sanitization approach. For experimental settings, we vary the privacy budget from 1 to 32 and apply the black-box instruction-based attack to LLaMA models. Figure 3, Figure 3, Figure 3 respectively provide the correlation between the Succ, Recall, Precision, and the privacy budget. We can observe that, for a greater privacy budget ($\epsilon\geq 8$), our attacks have higher precision and recall, indicating that our approach can attack the DP text sanitization approach under a practical privacy budget. In contrast, for a smaller privacy budget, both precision and recall are low, indicating that only a minimal amount of sensitive information can be recovered.

### v3

@@ VIII-C Large Language Models Used in the Experiments
VIII-D Ethical Considerations
Specifically, this paper discovered the capability of the modern large language models (LLMs) on recovering private content from differentially private prompts where the private information was sanitized. Such findings were based on experiments over the publicly available data and the open APIs to publicly available LLMs (see details in Section V). We faithfully followed the Terms of Service of LLMs when interacting with the LLMs. Since it was believed that differentially private outputs were not invertible, our work might have the following negative outcomes for public interests.
  • Disclosure. We discovered a vulnerability in the existing privacy-preserving technique of differential privacy. According to the paper, adversaries could use the latest implementations of LLMs to recover historical data that was intended to be protected by differential privacy. To mitigate this vulnerability, we discussed potential solutions in Section VII.
Overall, we did not engage in any malicious activities, such as exposing sensitive information, disrupting legitimate services, or causing financial or reputational harm to the LLM vendors providing these services.